OpenStudy (anonymous):

Find the area of a regular pentagon with a side length of 18 and an apothem of 25. 1075 1100 1125 1150 please show me the steps and answer

OpenStudy (anonymous):

The area can be found by multiplying half of the perimeter by the apothem. Here, the perimeter is 18*5=90 So, 25(45)=1125

jhonyy9 (jhonyy9):

area of one triangle will be 18*25/2 =9*25=225 so this we need multiplie by 5 because there is pentagon so 5*225=1125 - so the result right ,correct is 1125